OpenStudy (anonymous):

The line that passes through (1, 5) and (-2, 3) in standard form

OpenStudy (the_fizicx99):

"standard form" ? Do you mean Ax + By = C ? Anyways, so first find the slope, M = y2-y1/x2-x1 M = 3-5 / -2 - 1 M = -2 / -3 So your slope is 2/3 Now plug in y-y1=M(x-x1) y - 5 = 2/3(x-1) y - 5 = 2/3x - 0.666 +5 +5 y = 2/3x + 4.33 I don't know what you mean by standard form

OpenStudy (shamil98):

That is standard form @tHe_FiZiCx99 .

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3) that was the question I was given..

OpenStudy (shamil98):

Okay, first off. Find the slope of the two pairs of coordinates. \[slope = m = \frac{ y_2-y_1 }{ x_2-x_1 }\] \[m = \frac{ 3-5 }{ -2-1 } = \frac{ -2 }{ -3 } = \frac{ 2 }{ 3 }\] Then use the formula \[y - y_1 = m(x-x_1)\]

OpenStudy (shamil98):

\[y - 5 = \frac{ 2(x-1) }{ 3 }\] Multiply both sides by 3. \[3y - 15 = 2x - 2\] Change to standard form of \[ax +by = c\] \[-2x+ 3y = 13\]

OpenStudy (anonymous):

Find the slope of the line whose equation is 7y = 3x + 1

OpenStudy (shamil98):

7y = 3x +1 Remember the formula for slope-intercept is: \[y = mx + b\] where m = slope So, isolate y. \[\frac{ 7y }{ 7 } = \frac{ 3x }{ 7 } + \frac{ 3 }{ 7 }\] \[y = \frac{ 3 }{ 7 } x + \frac{ 3 }{ 7 }\] So slope = m = 3/7

OpenStudy (anonymous):

John made a rectangle pen for his dog using 28 feet of fencing. If the width of the pen is 2 feet more than one-half the length, what is the length and width of the pen? The length of the pen is

OpenStudy (shamil98):

Rectangle perimeter = 28 ft width = 2 + 1/2(L) Perimeter = 2L + 2W so 28 = 2L + 2(2+1/2(L)) 28 = 2L + 4 + L 24 = 3L 8 = L W = 2 + (1/2)L W = 2 + (1/2)8 W = 2 + 4 W = 6

OpenStudy (anonymous):

Find the slope of the line passing through the points (3, 8) and (-2, 5)

OpenStudy (shamil98):

Use the slope formula I posted above.

OpenStudy (anonymous):

I did.. I got .6 & that's not an answer on my choices..

OpenStudy (shamil98):

.6 = 6/10 = 3/5

OpenStudy (anonymous):

Oh, lol. I feel so dumb!

OpenStudy (anonymous):

2x+y=7 5x+y=9 Solve the system of equations

OpenStudy (shamil98):

y = -2x + 7 5x + (-2x +7) = 9 3x = 2 x= 2/3 y = -2(2/3) + 7 y = -4/3 + 7 y = -4/3 + 21/3 y = 17/3

OpenStudy (the_fizicx99):

( ._________________.)

OpenStudy (shamil98):

you noob.

OpenStudy (the_fizicx99):

Me? you did his hw =.= nub

OpenStudy (shamil98):

i got bored

OpenStudy (the_fizicx99):

o.0